The expression _r2 gives the area of a circle
with radius r. How many square meters of wheat were
destroyed when an alien ship made a crop circle of diameter
25 meters in the wheat field at the Southwind Ranch?
Round to the nearest tenth. Find _ on your calculator.

Answer by stanbon(75887)

Area = pi*r^2
r = 25
Area = pi*25^2
Area = 625*pi m^2
Area is approximately 1963.50 sq. meters
==========================================
Cheers,
Stan H.